A plane can fly 700 miles with the wind in 3.5 hours. the same plane travels 800 miles against the wind in 5 hours. what is the speed of the wind and the speed of the plane in still air????
Answer by checkley77(12844)

700/3.5=200 mph with the wind.
800/5=160 mph against the wind.
(200-160)/2=40/2=20 mph is the wind speed.
200-20=180 plane speed.
160+20=180 ditto.